studies indicate that the mean height of women is 63.7 with a standard deviations of 2.7 in.if a women is selected at random, what is the probability that she is 69 inches tall or taller
Answer #1

Solution :

Given that,

mean = = 63.7

standard deviation = = 2.7

P(x >69 ) = 1 - P(x<69 )

= 1 - P[(x -) / < (69-63.7) /2.7 ]

= 1 - P(z <1.96 )

Using z table

= 1 -  0.9750

probability= 0.0250